OpenStudy (anonymous):

solve using elimination 4y=-6x+30 7x+10y=35

OpenStudy (anonymous):

6x+4y=30 7x+10y=35

OpenStudy (anonymous):

-60x-40y=-300 28x+40=140

OpenStudy (anonymous):

-32x=-160 x=5

OpenStudy (anonymous):

plug x into any of the original equations

OpenStudy (anonymous):

what about y? do i need to do anything else?

OpenStudy (anonymous):

you plug in x to find y

OpenStudy (anonymous):

so plug in 5 where x is and solve for y

OpenStudy (anonymous):

nice thank you so much!!!!

OpenStudy (anonymous):

no problem :)